# arXiv:2403.09539v2[cs.CL] 15 Mar 2024

### v2

@@ 6.4 Basis-aware sampling
In a recent paper, Finlayson et al. (2023) propose a decoding algorithm that avoids type-I sampling errors by identifying tokens that must have non-zero true probability. Importantly, this method relies on knowing the basis of the LLM’s output space, and is therefore only available for LLMs whose image is known. Our proposed approach for finding the image of API-protected LLMs makes this decoding algorithm possible for API LLMs.
@@ 7 Mitigations
We consider a three proposals that LLM providers may take to guard against our methods. The first proposal is to remove API access to logprobs. This, however is theoretically insufficient, as Morris et al. (2023) show that it is possible to obtain full outputs using only information about the biased argmax token, albeit inefficiently in $v\log\epsilon$ API calls. Regardless of the theoretical result, providers may rely on the extreme inefficiency of the algorithm to protect the LLM. This appears to be the approach OpenAI took after learning about this vulnerability from Carlini et al. (2024), by always returning the top-$k$ unbiased logprobs. Our new proposed algorithm, however, brings the number of queries down to a more feasible $d\log\epsilon$ API calls once the initial work of finding the LLM image has finished, weakening the argument that the expensiveness of the algorithm is sufficient to disregard our technique.
@@
The next proposal is to remove API access to logit bias. This would be an effective way to protect the LLM, since there are no known methods to recover full outputs from such an API. However, the logit bias interface is presumably useful for many clients who might be disappointed by its shutdown.
@@
Lastly, we consider alternative LLM architectures that do not suffer from a softmax bottleneck. There are several such proposed architectures with good performance. Though this is the most expensive of the proposed defenses, due to the requirement of training a new LLM, it would have the beneficial side effect of also treating other tokenization issues that plague large-vocabulary LLMs (e.g., Itzhak and Levy, 2022). A transition to softmax-bottleneck-free LLMs would fully prevent our attack, since the model’s image would be the full output space.
8 Discussion
Overall, the potential applications of our methods can have great impact in building trust between the LLM API users and the providers; at the same time, none are particularly harmful to LLM providers. For instance, consider discovering the embedding size of the LLM, which is often non-public information in proprietary LLMs. However, access to this hyperparameter does not enable a competitor to fully recover the parameters of the LLM’s softmax matrix or boost performance of their own model; several other factors such as training data mix, human feedback and several other engineering choices are still critical and hidden. Even using model outputs to steal hidden prompts (see Section 6.3) is unlikely to have detrimental effects, as hidden prompt leakage is a known vulnerability and best practice dictates that no private information should be contained in these system prompts (Greshake et al., 2023; Liu et al., 2023). We find that the most dangerous consequence of our findings might simply be that model stealing methods that rely on full outputs get cheaper by a factor of $v/d$, which in the case of gpt-3.5-turbo amounts to about $25\times$.
On the other hand, allowing LLM API users to detect model changes builds trust between LLM providers and their customers, and leads to greater accountability and transparency for the providers. Our method can be used to implement efficient protocols for model auditing without exposing the model parameters or detailed configuration information, which may help model safety and privacy protection of personalized, proprietary models. We therefore take the position that our proposed methods and findings do not necessitate a change in LLM API best practices, but rather expand the tools available to API customers, while warning LLM providers of the information their APIs expose.
9 Simultaneous discovery
In a remarkable case of simultaneous discovery, Carlini et al. (2024) propose a very similar approach to ours for gaining insight into API-protected LLMs. Here we review a some interesting interactions between our work and theirs. First, we give an algorithm for obtaining full outputs in $v/k$ API calls, while their algorithm corresponds to our v/(k1)v/(k-1) algorithm. This has little impact on our final result, since our $v/k$ algorithm suffers from numerical instability, but it is an interesting theoretical result nonetheless. On the flip side, Carlini et al. propose an improved logprob-free method for getting full outputs from an API that takes advantage of parallel queries. This useful method is actually complementary to our $O(d)$ algorithm, since they can be combined to yield an even better algorithm for obtaining full outputs.
Next, our experiments with OpenAI’s API were fraught with issues of stochasticity: the API would return different results for the same query, leading us to develop the stochastically robust full output algorithm. Meanwhile, our colleagues did not appear to encounter such issues, perhaps because they had access to a more stable API endpoint than ours.
Lastly, Carlini et al. focus mostly on defenses and mitigations against attacks, while our own work focuses more on understanding the capabilities such attacks provide once an LLM image has been obtained. These approaches complement each other, since our work provides additional motivation for why (or why not) such defenses should be implemented.
@@ 10 Conclusion
In this paper we have laid out the fundamentals of how the softmax bottleneck imposes the low-rank constraints on LLM outputs, and shown how these constraints expose non-public information about API-protected LLMs. We have explained how this information, i.e., the LLM’s image, lends itself to a model signature, which has myriad uses—exposing model hyperparameters, recovering the output layer parameters, detecting specific types of model updates, and efficiently obtaining full outputs from the model, among others. We find that current protections against these attacks may be insufficient, and that the more effective guards we consider either eliminate sanctioned API use cases (i.e., logit bias), or are expensive to implement (i.e., changing model architecture). Lastly, we argue that this “vulnerability” may be more of a feature than a bug; since we believe that the benefits for API clients outweigh any known harms to LLM providers.
11 Acknowledgements
We thank Jack Morris and Vitaly Shmatikov for their insights and feedback on this work.
@@ A.1 Fast logprobs algorithm
$p_{i}=\frac{p^{\prime}_{i}}{\exp b_{\max}-\exp b_{\max}\sum_{j=1}^{k}p^{\prime}_{j}+\sum_{j=1}^{k}p^{\prime}_{j}}$ (6)
@@
$p^{\prime}_{i}=\frac{\exp(\ell_{i}+b_{\max})}{\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max})+\sum_{j=k+1}^{v}\exp\ell_{j}}$ (7)
@@
$\sum_{j=k+1}^{v}\exp\ell_{j}=\frac{\exp(\ell_{i}+b_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max}),$ (8)
@@
$\frac{\exp\ell_{i}}{p_{i}}-\sum_{j=1}^{k}\exp\ell_{j}=\frac{\exp(\ell_{i}+b_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max}).$ (9)
@@
$\frac{\exp\ell_{i}}{p_{i}}$ $=\frac{\exp(\ell_{i}+b_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max})+\sum_{j=1}^{k}\exp\ell_{j}$ (10)
$p_{i}$ $=\frac{\exp\ell_{i}}{\frac{\exp(\ell_{i}+b_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max})+\sum_{j=1}^{k}\exp\ell_{j}}$ (11)
@@
$p_{i}$ $=\frac{p^{\prime}_{i}\exp\ell_{i}}{\exp(\ell_{i}+b_{\max})-p^{\prime}_{i}\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max})+p^{\prime}_{i}\sum_{j=1}^{k}\exp\ell_{j}}$ (12)
$=\frac{p^{\prime 2}_{i}\exp(-b_{\max})\exp(\ell_{i}+b_{\max})}{\exp(\ell_{i}+b_{\max})-p^{\prime}_{i}\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max})+p^{\prime}_{i}\exp(-b_{\max})\sum_{j=1}^{k}\exp(\ell_{j}+b_{\max})}$ (13)
@@
$p_{i}$ $=\frac{p^{\prime 2}_{i}\exp(-b_{\max})}{p^{\prime}_{i}-p^{\prime}_{i}\sum_{j=1}^{k}p^{\prime}_{j}+p^{\prime}_{i}\exp(-b_{\max})\sum_{j=1}^{k}p^{\prime}_{j}}$ (14)
$=\frac{p^{\prime}_{i}\exp(-b_{\max})}{1-\sum_{j=1}^{k}p^{\prime}_{j}+\exp(-b_{\max})\sum_{j=1}^{k}p^{\prime}_{j}}$ (15)
$=\frac{p^{\prime}_{i}}{\exp b_{\max}-\exp b_{\max}\sum_{j=1}^{k}p^{\prime}_{j}+\sum_{j=1}^{k}p^{\prime}_{j},}$ (16)
@@ A.2 Numerically stable algorithm
Combining these agian, we get
@@
Next, we use the fact that $\exp\ell^{\prime}_{i}=\exp b_{\max}\exp\ell$ to get
$\frac{p_{v}^{\prime}p_{i}}{p_{i}^{\prime}p_{v}}=\exp(-b_{\max})\,.$ (22)
@@ A.3 Stochastically robust algorithm
Using our result from Appendix A.2, we have
$p_{i}$ $=\exp(\log p^{\prime}_{i}-b_{\max}-\log p^{\prime}_{v}+\log p_{v})$ (25)
$p_{i}$ $=\exp(\log p^{\prime}_{i}-b_{\max}-\log p^{\prime}_{v-1}+\log p_{v-1}).$ (26)

### v3

@@ 7.4 Basis-aware sampling
In a recent paper, Finlayson et al. (2024) propose a decoding algorithm that avoids type-I sampling errors by identifying tokens that must have non-zero true probability. Importantly, this method relies on knowing the basis of the LLM’s output space, and is therefore only available for LLMs whose image is known. Our approach for finding the model image makes this decoding algorithm possible for API LLMs.
@@ 8 Mitigations
We consider three proposals that LLM providers may take to guard against our methods. The first proposal is to limit or entirely remove API access to logprobs. This is theoretically insufficient, as Morris et al. (2023) show that it is possible to obtain full outputs using only information about the biased argmax token, albeit inefficiently in $v\log(\beta/\epsilon)$ API calls. Providers may rely on the extreme inefficiency of the algorithm to protect the LLM, as OpenAI appeared to do in response to Carlini et al. (2024). However, our algorithm in §4 brings the cost down to a feasible $O(d\log(\beta/\epsilon))$ API calls per output once the initial work of finding the LLM image finishes, the result of which can be made public and reused.
@@
The second proposal is to remove API access to logit bias. This would be effective, since there are no known methods to recover full outputs from such an API. However, logit bias has several useful applications, such as for blocking undesirable tokens and controlling text generation, and making such a restriction could seriously degrade the usefulness of the API.
@@
Lastly, we consider alternative LLM architectures that do not suffer from a softmax bottleneck. There are several such proposed architectures with good performance (e.g., Xue et al., 2021; Yu et al., 2023; Wang et al., 2024). Though this is the most expensive of the proposed defenses, due to the requirement of training a new LLM, it would have the beneficial side effect of also treating other tokenization issues that plague large-vocabulary LLMs (e.g., Itzhak and Levy, 2022). A transition to softmax-bottleneck-free LLMs would fully prevent our attack, since the model’s image would be the full output space.
@@ 9 Conclusion
We have shown how the low-rank constraints imposed by the softmax bottleneck expose non-public information about API-protected LLMs. We also demonstrated how this information can be used to efficiently extract full model outputs, expose model hyperparameters, function as a unique model signature for auditing purposes, and even detect specific types of model updates, among other uses. We find that some current protections against these attacks are insufficient, and that the more effective guards tend to inhibit sanctioned API use cases (i.e., logit bias), or are expensive to implement (i.e., changing model architecture).
Overall, we find that the benefits of our proposed methods outweigh the harms to LLM providers. For instance, allowing LLM API users to detect model changes builds trust between LLM providers and their customers, and leads to greater accountability and transparency for the providers. Our method can also be used to implement efficient protocols for model auditing without exposing the model weights. On the other hand, discovering the embedding size of the LLM does not enable a competitor to fully recover the parameters of the LLM’s softmax matrix or boost performance of their own model. Even efficiently extracting full outputs for model stealing or inversion (§4) is unlikely to have detrimental effects, as these are already known threats and are only made more efficient by our methods. We therefore believe that our proposed methods and findings do not necessitate a change in LLM API best practices, but rather expand the tools available to API customers, while informing LLM providers of what information their APIs expose.
@@ A.1 Fast logprobs algorithm
$p_{i}=\frac{p^{\prime}_{i}}{\exp\beta_{\max}-\exp\beta_{\max}\sum_{j=1}^{k}p^{\prime}_{j}+\sum_{j=1}^{k}p^{\prime}_{j}}$ (4)
@@
$p^{\prime}_{i}=\frac{\exp(\ell_{i}+\beta_{\max})}{\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max})+\sum_{j=k+1}^{v}\exp\ell_{j}}$ (5)
@@
$\sum_{j=k+1}^{v}\exp\ell_{j}=\frac{\exp(\ell_{i}+\beta_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max}),$ (6)
@@
$\frac{\exp\ell_{i}}{p_{i}}-\sum_{j=1}^{k}\exp\ell_{j}=\frac{\exp(\ell_{i}+\beta_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max}).$ (7)
@@
$\frac{\exp\ell_{i}}{p_{i}}$ $=\frac{\exp(\ell_{i}+\beta_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max})+\sum_{j=1}^{k}\exp\ell_{j}$ (8)
$p_{i}$ $=\frac{\exp\ell_{i}}{\frac{\exp(\ell_{i}+\beta_{\max})}{p^{\prime}_{i}}-\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max})+\sum_{j=1}^{k}\exp\ell_{j}}$ (9)
@@
$p_{i}$ $=\frac{p^{\prime}_{i}\exp\ell_{i}}{\exp(\ell_{i}+\beta_{\max})-p^{\prime}_{i}\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max})+p^{\prime}_{i}\sum_{j=1}^{k}\exp\ell_{j}}$ (10)
$=\frac{p^{\prime 2}_{i}\exp(-\beta_{\max})\exp(\ell_{i}+\beta_{\max})}{\exp(\ell_{i}+\beta_{\max})-p^{\prime}_{i}\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max})+p^{\prime}_{i}\exp(-\beta_{\max})\sum_{j=1}^{k}\exp(\ell_{j}+\beta_{\max})}$ (11)
@@
$p_{i}$ $=\frac{p^{\prime 2}_{i}\exp(-\beta_{\max})}{p^{\prime}_{i}-p^{\prime}_{i}\sum_{j=1}^{k}p^{\prime}_{j}+p^{\prime}_{i}\exp(-\beta_{\max})\sum_{j=1}^{k}p^{\prime}_{j}}$ (13)
$=\frac{p^{\prime}_{i}\exp(-\beta_{\max})}{1-\sum_{j=1}^{k}p^{\prime}_{j}+\exp(-\beta_{\max})\sum_{j=1}^{k}p^{\prime}_{j}}$ (14)
$=\frac{p^{\prime}_{i}}{\exp\beta_{\max}-\exp\beta_{\max}\sum_{j=1}^{k}p^{\prime}_{j}+\sum_{j=1}^{k}p^{\prime}_{j},}$ (15)
@@ A.2 Numerically stable algorithm
Combining these again, we get
@@
Next, we use the fact that $\exp\ell^{\prime}_{i}=\exp\beta_{\max}\exp\ell$ to get
$\frac{p_{v}^{\prime}p_{i}}{p_{i}^{\prime}p_{v}}=\exp(-\beta_{\max})\,.$ (21)
@@ A.3 Stochastically robust algorithm
Using our result from §A.2, we have
$p_{i}$ $=\exp(\log p^{\prime}_{i}-\beta_{\max}-\log p^{\prime}_{v}+\log p_{v})$ (24)
$p_{i}$ $=\exp(\log p^{\prime}_{i}-\beta_{\max}-\log p^{\prime}_{v-1}+\log p_{v-1}).$ (25)
@@
Appendix B Solving for full outputs using the LLM image
Given a matrix $\boldsymbol{P}\in\Delta_{v}^{d}$ whose columns span the LLM image, and $p_{1},p_{2},\ldots,p_{d}$, we can solve for $\boldsymbol{p}\in\Delta_{v}$. To do this, we will use the additive log ratio ($\operatorname{alr}$) transform, which is an isomorphism $\Delta_{v}\to\mathbb{R}^{v-1}$ and is defined as
$\operatorname{alr}(\boldsymbol{p})=\left(\log\frac{p_{2}}{p_{1}},\log\frac{p_{3}}{p_{1}},\ldots,\log\frac{p_{v}}{p_{1}}\right)$ (26)
$\operatorname{alr}(\boldsymbol{P})=\begin{bmatrix}\operatorname{alr}(\boldsymbol{p}^{1})&\operatorname{alr}(\boldsymbol{p}^{2})&\cdots&\operatorname{alr}(\boldsymbol{p}^{d})\end{bmatrix}\in\mathbb{R}^{(v-1)\times d}$ (27)
$\operatorname{alr}^{-1}(\boldsymbol{x})=\frac{1}{1+\sum_{i=1}^{v-1}\exp x_{i}}\cdot\left(1,\exp x_{1},\exp x_{2},\ldots,\exp x_{v-1}\right).$ (29)